ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 149, 2005

Host, B. and B. Kra Convergence of polynomial ergodic averages  1

Hubička, J. and J. Nešetřil Finite presentation of homogeneous graphs, posets and Ramsey classes

21

Kaimanovich, V. A. Amenability and the Liouville property  45
Kamae, T. Numeration systems, fractals and stochastic processes  87

Kleinbock, D. and B. Weiss Badly approximable vectors on fractals

137

Knieper, G. The uniqueness of the maximal measure for geodesic flows on symmetric spaces of higher rank

171

Kra, B. See B. Host

Levin, G. and F. Przytycki On Hausdorff dimension of some Cantor attractors

185

Lindenstrauss, E. Rigidity of multiparameter actions  199

Lindenstrauss, E. and K. Schmidt Symbolic representations of nonexpansive group automorphisms

227

Lubotzky, A., B. Samuels and U. Vishne Ramanujan complexes

of type Ã_{d} 267

Nešetřil, J. See J. Hubička

Oh, H. The Ruziewicz problem and distributing points on homogeneous spaces of a compact Lie group

301

Przytycki, F. See G. Levin

Samuels, B. See A. Lubotzky

Schmidt, K. See E. Lindenstrauss

Thuswaldner, J. M. and R. F. Tichy Waring’s problem with digital restrictions

317

Tichy, R. F. See J. M. Thuswaldner

Vishne, U. See A. Lubotzky

Weiss, B. Some remarks on filtering and prediction of stationary processes

345

Weiss, B. See D. Kleinbock